OpenStudy (anonymous):

Kim knows she should keep her fat intake to less than 130 grams per day if on a 4,000 calorie diet. She is on a 1,500 calorie per day diet. How many grams of fat should she limit herself to each day?

OpenStudy (phi):

do you know about ratios and proportions ?

OpenStudy (phi):

write a ratio of fat grams/ calories

OpenStudy (anonymous):

yes but it has been I really long time since I have done stuff like this

OpenStudy (phi):

Use this info to write the first ratio: 130 grams per day if on a 4,000 calorie diet.

OpenStudy (anonymous):

130/4000

OpenStudy (phi):

and for the second ratio, call the "unknown fat grams" x with 1500 calories

OpenStudy (anonymous):

x/1500

OpenStudy (phi):

and set the ratios equal \[ \frac{x}{1500} = \frac{130}{4000} \] to "solve for x" multiply both sides by 1500, like this: \[ \frac{x}{1500} \cdot 1500= \frac{130}{4000} \cdot 1500\] and simplify both sides

OpenStudy (anonymous):

I got 195000

OpenStudy (phi):

on the left side x *1500/1500 is x*1 = x on the right side, 130*1500/4000 try again

OpenStudy (phi):

apparently you did 130*1500 but did not divide by 4000

OpenStudy (anonymous):

oh that would be 48.75

OpenStudy (phi):

yes. so the answer is 48.75 grams, or if you round to the nearest gram, about 49 grams